# ToddlerBot: Open-Source ML-Compatible Humanoid Platform for Loco-Manipulation

### v2

@@ III-B Capability - Design Objective
Mechatronic Design. As shown in Figure 3, ToddlerBot replicates the structure of an adult human body to maximize the number of active DoFs, enabling human-like motion and versatile loco-manipulation tasks. Each arm features seven DoFs with spur gears for axis-aligned transmission, ensuring high functionality and reachability. Each leg has six DoFs: three at the hip for a wide range of motion and optimized walking strides, a parallel linkage at the knee to reduce inertia, and two at the ankle for stable locomotion. The neck, with two DoFs, incorporates parallel linkages at the pitch joint for a compact design, allowing expressive motion and full head mobility. The waist, also with two DoFs, uses coupled bevel gears to balance the space budget and effectively transmit power from two motors for yaw and roll actuation, which enables whole-body control. Additionally, ToddlerBot offers two end-effector designs: a parallel jaw gripper for grasping and a compliant palm for tasks requiring a palm-like posture. The end-effector designs can be switched quickly within two minutes by removing a few screws.
For each active DoF, we maximize the range of motion by optimizing geometries to prevent self-collisions as shown in Section VIII-C of the supplementary material. To address additional concerns such as space constraints, axis alignment, and inertia reduction, we integrate three primary transmission mechanisms in the mechanical design: spur gears, coupled bevel gears, and parallel linkages, as discussed in Section VIII-D of the supplementary material.
Motor Selection. We choose Dynamixel motors because of their robustness, reliability, and accessibility. Different types of Dynamixel motors were selected for various joints based on space constraints, torque requirements, and cost considerations. In terms of communication speed, Dynamixel motors communicate via a $5\mathrm{V}$ TTL protocol running at 2M baudrate, providing full-state feedback for all 30 motors at $50\leavevmode\nobreak\ \mathrm{Hz}$ using an off-the-shelf communication board. More detailed considerations are discussed in Section VIII-E of the supplementary material.
@@
Sensors, Compute, and Power. To enhance ToddlerBot’s capabilities, we integrate a comprehensive set of sensors and computational components. Two fisheye cameras are included to expand the field of view, improving situational awareness. An Inertial Measurement Unit (IMU) is placed in the chest to provide state feedback for precise control. A speaker and microphones facilitate seamless communication with humans and other ToddlerBot instances. We provide a conversation example between two ToddlerBot instances in the supplementary video. The onboard computation is powered by a Jetson Orin NX 16GB, enabling real-time inference of machine learning models. Power management is handled by a custom-designed power distribution board with dual $12\mathrm{V}$ converters. This system efficiently regulates the $12\mathrm{V}$ rail and distributes power to various components, meeting the high energy demands while fitting within the limited space of the torso. The power budget is discussed in Section VIII-F of the supplementary material.
Figure 4: Zero-point Calibration. We 3D-print devices for the plug-and-play zero-point calibration procedure: orange for the arm, yellow for the neck, red for the hip, and beige for the ankle. Arrows indicate the insertion direction, and the zero-point is fixed once the devices click into place.
III-C ML-Compatibility - Design Objective
Digital Twin. A high-fidelity digital twin is essential for high-quality simulation data collection and thus zero-shot sim-to-real transfer. We divide the digital twin development into two key components: zero-point calibration for correct kinematics and motor system identification for accurate dynamics.
Since Dynamixel motors lack an inherent zero point, a reliable method is needed to recalibrate after assembly, which is frequent during repairs or design iterations. As shown in Figure 4, we design calibration devices in CAD that quickly align the robot to its zero point, defined as standing with both arms besides the body. The process takes less than a minute.
With kinematics correctly calibrated, the next challenge is accurately identifying dynamics parameters. Firstly, while it may seem trivial, it is essential to carefully weigh all components to ensure an accurate weight distribution for the robot. Inspired by Haarnoja et al. [17], we collect sysID data by commanding the motors to track a chirp signal and use the resulting position tracking data to fit an actuation model as described in Grandia et al. [13]. The actuation model is detailed in Section VIII-G of the supplementary material. We assume that motors of the same model have nearly identical dynamics parameters due to manufacturing consistency. This assumption is empirically validated by performing sysID only five times—once for each Dynamixel motor model—and successfully transferring policies to a second ToddlerBot instance without additional sysID.
@@ III-C ML-Compatibility - Design Objective
Teleoperation Device. Previous works [1, 48] demonstrate that teleoperation is one of the most effective ways to gather real-world data. Inspired by them, we develop a second upper body of ToddlerBot as the leader arms. Two force-sensitive resistors (FSRs) are embedded in the gripping area of the end effectors to detect compression force from the operator, allowing gripper movement based on force input.
We use a handheld gaming computer (either Steam Deck or ROG Ally X) to control the other body parts. The joysticks send velocity commands to walk, turn, and squat. Buttons trigger either programmed or trained policies, where holding a button executes the policy continuously, and releasing it stops the action. Buttons also provide direct control over neck and waist movements. The detailed mapping is described in Section VIII-H of the supplementary material.
IV System Control
IV-A Keyframe Animation
@@
Keyframe animation is a cornerstone of character animation, but it provides only kinematic data, with no guarantee of dynamic feasibility [21]. To address this, we developed software integrating MuJoCo [45] with a GUI, enabling real-time tuning and validation of keyframes and motion trajectories generated through linear interpolation with user-defined timings. With our high-fidelity digital twin, we efficiently generate open-loop trajectories such as cuddling, push-ups, and pull-ups that can be executed zero-shot in the real world.
IV-B Reinforcement Learning
@@
For walking and turning, we train a reinforcement learning (RL) policy, $\pi(\bm{\mathrm{a}}_{t}|\bm{\mathrm{s}}_{t})$, which outputs $\bm{\mathrm{a}}_{t}$ as joint position setpoints for proportional-derivative (PD) controllers, based on the observable state $\bm{\mathrm{s}}_{t}$:
@@
where $\bm{\phi}_{t}$ is a phase signal, $\bm{c}_{t}$ represents velocity commands, $\Delta\bm{q}_{t}$ denotes the position offset relative to the neutral pose $\bm{q}_{0}$, $\bm{a}_{t-1}$ is the action from the previous time step, $\bm{\theta}_{t}$ represents the torso orientation, and $\bm{\omega}_{t}$ is the torso’s angular velocity. All inputs are normalized to ensure stable learning.
During PPO policy training [40], the environment generates the next state, $\bm{s}_{t+1}$, updates the phase signal, and returns a scalar reward $\mathrm{r}_{t}=\mathrm{r}(\bm{s}_{t},\bm{a}_{t},\bm{s}_{t+1},\bm{\phi}_{t},\bm{c}_{t})$. Following standard practice [13], the reward is decomposed as:
@@
Among these components, $\mathrm{r}_{t}^{\text{imitation}}$ encourages accurate imitation of the reference walking motion, which is generated using a closed-form ZMP (Zero Moment Point) solution [43]. $\mathrm{r}_{t}^{\text{regularization}}$ incorporates heuristics of ideal walking motion, penalizes joint torques, and promotes smooth actions to minimize unnecessary movements. A survival reward $\mathrm{r}_{t}^{\text{survival}}$ prevents early episode termination during training. Additional details are provided in Section VIII-I of the supplementary material.
Figure 5: Arm Span and Payload. We show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, about 14 times the torso size. Additionally, ToddlerBot can lift weights up to $1484\leavevmode\nobreak\ \mathrm{g}$, which is 40% of its body weight ($3484\leavevmode\nobreak\ \mathrm{g}$).
IV-C Imitation Learning
Real-world data collection involves a human operator guiding the leader’s arms to teleoperate the follower’s arms while using a joystick and buttons on a remote controller to control body movements.
@@ IV-C Imitation Learning
During data collection, when the upper body tracks the position commands from the leader arms, ToddlerBot’s lower body employs a two-layer PD controller to maintain balance. The first layer is a Center of Mass (CoM) PD controller, which keeps the CoM close to the center of the support polygon. The first layer addresses CoM shifts caused by arm movements, while the second layer manages lifting heavy objects. The second layer is a torso pitch PD controller, which uses IMU readings to ensure the torso remains upright.
@@
Figure 6: Trajectory Tracking. We present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile. Both raw and smoothed linear and angular velocity tracking are displayed, with real-world results averaged across trials.
Table II: Tracking Errors in Simulation versus the Real World
Tracking Errors Simulation Real-World
Position [m] 0.082 0.133 $\pm$ 0.018
Linear Velocity [m/s] 0.016 0.032 $\pm$ 0.002
Angular Velocity [rad/s] 0.056 0.113 $\pm$ 0.010
With this setup, we can collect 60 trajectories in just 20 minutes for both bimanual and full-body manipulation tasks. The motor positions of the leader’s arms are recorded as the actions, while the motor positions of the follower robot, along with the RGB images captured from its camera, are recorded as observations. This data is subsequently used to train a diffusion policy [3], with further details provided in Section VIII-J of the supplementary material.
V Experiments
We conducted a series of experiments to show that ToddlerBot is capable, ML-compatible, and reproducible.
Figure 7: Experiment Results. We present a series of photos showcasing four different tasks: push-up, pull-up, bimanual, and full-body manipulation, demonstrating ToddlerBot’s capability in challenging locomotion and manipulation tasks.
V-A Capability: Arm Span, Payload, and Endurance
@@
To evaluate ToddlerBot’s arm span and payload capacity, we teleoperate it to hug a test object using the compliant palm gripper while maintaining balance. In the arm span test, we show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
@@
The payload test assesses both the upper body’s lifting capacity and the lower body’s ability to maintain balance. ToddlerBot successfully lifts up to $1484\leavevmode\nobreak\ \mathrm{g}$, 40% of its total weight ($3484\leavevmode\nobreak\ \mathrm{g}$). To eliminate friction effects, we use a 3D-printed cup for the compliant palm gripper to lock in securely. To determine the true limit, we incrementally add screws to the cup until ToddlerBot falls over. Results are shown in Figure 5.
@@
In the endurance test, ToddlerBot starts with a fully charged battery, running the walking RL policy while stepping in place. ToddlerBot achieved the longest streak of 19 minutes without falling. Over time, increased motor temperatures gradually pushed it outside the policy’s training distribution, leading to more frequent falls. ToddlerBot withstands up to 7 falls before breaking, but even then, repairs are quick—requiring only 21 minutes of 3D printing and 14 minutes of assembly, including removing the damaged part, installing the replacement, performing zero-point calibration, and rerunning the script.
V-B Capability: Push-ups and Pull-ups
@@
To demonstrate ToddlerBot’s ability to execute expressive and dynamic motions, we program push-ups and pull-ups in our keyframe software and perform a zero-shot sim-to-real transfer. For pull-ups, we use an AprilTag to help ToddlerBot accurately locate the horizontal bar. Both tasks require strong limbs, balanced upper-lower body strength, and precise coordination, particularly when ToddlerBot transitions from a planking pose to standing after push-ups and when it releases the horizontal bar and lands after pull-ups. These open-loop transfers require only a single motion trajectory designed in simulation, highlighting the fidelity of our digital twin. Figure 7 showcases both tasks through a series of images.
Figure 8: Skill Chaining. We demonstrate that ToddlerBot can seamlessly transition from a DP-based grasping skill to an RL-trained walking policy, enabling it to grasp the wagon handle and push it forward while maintaining its grip.
@@
Figure 9: Two-instance Collaboration. In this task, two instances of ToddlerBot, Arya and Toddy, collaborate to clean up a toy session. (1) The scene begins with a pink octopus on the table, a purple octopus on the ground, a wagon, and Arya standing next to the table. (2) Arya picks up the pink octopus from the table and places it in the wagon. (3) Arya walks to the wagon handle. (4) Arya runs a grasping policy to grasp the wagon handle, while Toddy walks over. (5) Arya pushes the wagon toward the purple octopus and waits for Toddy. (6) Toddy reaches the pickup position. (7) Toddy kneels and picks up the purple octopus. (8) Finally, Arya and Toddy walk side by side, leaving the scene together.
V-C ML-Compatibility: Omnidirectional Walking
@@
To demonstrate ToddlerBot’s locomotion capabilities, we train RL walking policies to follow a square trajectory with a predefined velocity profile. Figure 6 and Table II present the results, with real-world tracking data collected via motion capture. Due to RL policy limitations, both simulation and real-world tracking deviate from the command, primarily because the learned walking policy struggles with in-place rotation, causing translation offsets. However, the sim-to-real gap is notably smaller than the tracking gap, supporting a successful zero-shot transfer. Additionally, we report a position tracking error variance of $0.018\leavevmode\nobreak\ \mathrm{m}$, linear velocity tracking error of $0.002\leavevmode\nobreak\ \mathrm{m/s}$, and angular velocity tracking error of $0.01\leavevmode\nobreak\ \mathrm{rad/s}$, which demonstrates good repeatability.
V-D ML-Compatibility: Vision-based Manipulation
@@
We show ToddlerBot’s ability to perform bimanual manipulation and full-body manipulation by transferring octopus toys from a table and the ground to a wagon. Both tasks are trained with RGB-based diffusion policy [3] with 60 demonstrations. Across 20 test trials, we achieve a 90% success rate for bimanual manipulation and 75% for full-body manipulation. We leverage a combination of open-loop motions and closed-loop policies to enhance data collection efficiency. In the bimanual task, the torso rotating and releasing motions are open-loop, while in full-body manipulation, kneeling down is open-loop. The same motions are also used during policy evaluation. ToddlerBot’s onboard computing runs a 300M-parameter diffusion policy with about $100\leavevmode\nobreak\ \mathrm{ms}$ latency, enabling real-time operation. Results are shown in Figure 7.
V-E ML-Compatibility: Skill Chaining
@@
To test the system’s ability to combine locomotion and manipulation skills, we test the wagon-pushing task (Figure 8), which requires the robot to first perform vision-based grasping and then walk forward. To do so, ToddlerBot first executes a diffusion policy to grasp the handle, while maintaining that pose, switched to the RL policy to push the wagon forward. To enable ToddlerBot to walk while maintaining its grip, we sample the end pose of the upper body from the grasping policy’s training data (60 demonstrations) during RL training.
V-F Reproducibility: Hardware and Policies
To demonstrate hardware reproducibility, we recruit a CS-major student with no prior hardware experience to assemble a second ToddlerBot using our manuals as shown in Figure 15 of the supplementary material and our open-source assembly videos. The student independently completes the assembly in three days, including the time required for 3D printing.
@@ V-F Reproducibility: Hardware and Policies
For policy reproducibility, we run the manipulation policy trained on data collected with one instance on the other instance, achieving the same success rate of 90% across 20 trials. We also successfully transfer the RL walking policy between both robots. To further showcase the equivalent performance of both ToddlerBot instances, we have them collaborate on a long-horizon room tidying task, as shown in Figure 9.
VI Limitations and Future Work
ToddlerBot’s performance in more agile tasks is constrained by the off-the-shelf motors’ max speed, max torque, and communication speed. Rather than achieving superhuman capabilities, ToddlerBot aligns more closely with average human performance in loco-manipulation tasks. Additionally, its scale limits interaction with human-sized objects, though this does not hinder research if appropriately sized objects are used.
To overcome these limitations, we are developing customized communication boards and improving motor system identification to maximize performance. We also aim to improve sensing capabilities, including stereo vision for depth perception, additional IMUs for improved state estimation, and tactile sensors for richer feedback.
VII Conclusions

### v3

@@ 3.2 Capability - Design Objective
Mechatronic Design. As shown in Figure 2, ToddlerBot replicates the structure of an adult human body to maximize the number of active DoFs, enabling human-like motion and versatile loco-manipulation tasks. Each arm features seven DoFs with spur gears for axis-aligned transmission, ensuring high functionality and reachability. Each leg has six DoFs: three at the hip for a wide range of motion and optimized walking strides, a parallel linkage at the knee to reduce inertia, and two at the ankle for stable locomotion. The neck, with two DoFs, incorporates parallel linkages at the pitch joint for a compact design, allowing expressive motion and full head mobility. The waist, also with two DoFs, uses coupled bevel gears to balance the space budget and effectively transmit power from two motors for yaw and roll actuation, which enables whole-body control. Additionally, ToddlerBot offers two end-effector designs: a parallel jaw gripper for grasping and a compliant palm for tasks requiring a palm-like posture. The end-effector designs can be switched quickly within two minutes. For each active DoF, we maximize the range of motion by optimizing geometries to prevent self-collisions (Appendix 8.3). To address concerns such as space constraints, axis alignment, and inertia reduction, we integrated a combination of three transmission mechanisms in the design: spur gears, coupled bevel gears, and parallel linkages (Appendix 8.4).
@@
Sensors, Compute, and Power. To enhance ToddlerBot’s capabilities, we integrate a comprehensive set of sensors and computational components. Two fisheye cameras are included to expand the field of view. An Inertial Measurement Unit (IMU) is placed in the chest to provide state feedback. A speaker and microphones facilitate communication with humans and other robots. We provide a conversation example between two ToddlerBot instances in the supplementary video. The onboard computation is powered by a Jetson Orin NX 16GB, enabling real-time inference of ML models. Based on our proposed power metric (Appendix 8.5), we choose Dynamixel motors for their reliability and accessibility (Appendix 8.6). Power management is handled by a custom-designed power distribution board which can be easily ordered online (Appendix 8.7).
3.3 ML-Compatibility - Design Objective
Digital Twin. A high-fidelity digital twin is fundamental for high-quality simulation data collection and zero-shot sim-to-real transfer. We divide the digital twin development into two key components: zero-point calibration for correct kinematics and motor system identification for accurate dynamics. Since Dynamixel motors lack an absolute zero point, a reliable method is needed to calibrate after reassembly. We designed a set of calibration devices that can align the robot to the desired zero point within a minute, defined as standing with both arms parallel to the body (Appendix 8.8). For accurate sysID, we designed a motor test bed and automatic test procedures described in Appendix 8.9 to measure the range of dynamics parameters. Next, we collect actuator position tracking data by commanding a chirp signal [27], and use the position tracking result to fit an actuation model [42] with optimization, as detailed in Appendix 8.10. Testing across instances shows Dynamixel motors of the same model have nearly identical dynamics parameters. This is also validated by successful direct deployment of policies to a second ToddlerBot instance without additional sysID.
@@ 3.3 ML-Compatibility - Design Objective
Teleoperation Device. Inspired by previous works [4, 43], we develop a second upper body of ToddlerBot as the leader arms to collect high-quality real-world data. Two force-sensitive resistors (FSRs) are embedded in the gripping area of the end effectors to detect compression force from the operator, allowing gripper movement based on force input. We use a handheld gaming computer (either Steam Deck or ROG Ally X) to control the other body parts. The joysticks send velocity commands to walk, turn, and squat. Buttons trigger either programmed or trained policies and provide direct control over neck and waist movements. The detailed mapping is described in Appendix 8.11.
4 System Control
4.1 Keyframe Animation
@@
Keyframe animation is widely used in character animation, but it provides only kinematics data, lacking a guarantee of dynamic feasibility [44]. To address this, we develop a tool integrating MuJoCo [45] with a GUI, enabling real-time tuning and validation of keyframe motion trajectories. Combined with high-fidelity digital twin, we can efficiently generate open-loop trajectories such as cuddling, push-ups, and pull-ups that can be executed zero-shot in the real world.
4.2 Reinforcement Learning
@@
For walking, we train a reinforcement learning (RL) policy, $\pi(\bm{\mathrm{a}}_{t}|\bm{\mathrm{s}}_{t})$, which outputs $\bm{\mathrm{a}}_{t}$ as joint position setpoints for proportional-derivative (PD) controllers, based on the observable state $\bm{\mathrm{s}}_{t}$:
@@
where $\bm{\phi}_{t}$ is a phase signal, $\bm{c}_{t}$ represents velocity commands, $\Delta\bm{q}_{t}$ denotes the position offset relative to the neutral pose $\bm{q}_{0}$, $\bm{a}_{t-1}$ is the action from the previous time step, $\bm{\theta}_{t}$ represents the torso orientation, and $\bm{\omega}_{t}$ is the torso’s angular velocity. During PPO policy training [46], the environment generates the next state, $\bm{s}_{t+1}$, updates the phase signal, and returns a scalar reward $\mathrm{r}_{t}=\mathrm{r}(\bm{s}_{t},\bm{a}_{t},\bm{s}_{t+1},\bm{\phi}_{t},\bm{c}_{t})$. Following standard practice [42], the reward is decomposed as:
@@
Among these components, $\mathrm{r}_{t}^{\text{imitation}}$ encourages accurate imitation of the reference walking motion, which is generated using a closed-form ZMP (Zero Moment Point) solution [47]. $\mathrm{r}_{t}^{\text{regularization}}$ incorporates heuristics of ideal walking motion, penalizes joint torques, and promotes smooth actions to minimize unnecessary movements. A survival reward $\mathrm{r}_{t}^{\text{survival}}$ prevents early episode termination during training. Additional details are provided in Appendix 8.12.
4.3 Imitation Learning
@@
Real-world data collection involves a human operator guiding the leader’s arms to teleoperate the follower’s arms while using a joystick and buttons on a handheld game controller to control body movements. During data collection, when the upper body tracks the position commands from the leader arms, ToddlerBot’s lower body employs a two-layer PD controller to actively maintain balance. The first layer is a Center of Mass (CoM) PD controller, which keeps the CoM close to the center of the support polygon. The second layer is a torso pitch PD controller, which uses IMU readings to ensure the torso remains upright. The first layer addresses CoM shifts caused by arm movements, while the second layer compensates when lifting heavy objects. With this setup, we can collect 60 trajectories in just 20 minutes for both bimanual and full-body manipulation tasks. The motor positions of the leader’s arms are recorded as the actions, while the motor positions of the follower robot, along with the RGB images captured from its camera, are recorded as observations. We train a diffusion policy [48] on this dataset, with further details provided in Appendix 8.13.
5 Experiments
@@
Figure 3: Arm Span, Payload, and Trajectory Tracking. On the left, we show that with a torso dimension of $13\times 9\times 12\leavevmode\nobreak\ \mathrm{cm}^{3}$, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, about 14 times the torso size. Additionally, ToddlerBot can lift weights up to $1484\leavevmode\nobreak\ \mathrm{g}$, which is 40% of its body weight ($3484\leavevmode\nobreak\ \mathrm{g}$). On the right, we present ten consecutive real-world rollouts of an RL walking policy tracking a square trajectory with a predefined velocity profile. Both raw and smoothed linear and angular velocity tracking are displayed, with real-world results averaged across trials.
@@
Capability: Arm Span, Payload, and Endurance. To evaluate ToddlerBot’s arm span, we teleoperate it to hug a large object using the compliant palm gripper while maintaining balance. We show that, ToddlerBot can grasp objects up to $27\times 24\times 31\leavevmode\nobreak\ \mathrm{cm}^{3}$, approximately 14 times its torso volume.
@@
The payload test assesses both the upper body’s lifting capacity and the lower body’s ability to maintain balance. ToddlerBot successfully lifts up to $1484\leavevmode\nobreak\ \mathrm{g}$, 40% of its total weight ($3484\leavevmode\nobreak\ \mathrm{g}$). To eliminate friction effects, we use a 3D-printed cup for the gripper to lock in securely. We incrementally add screws to the cup until ToddlerBot falls over. Results are shown on the left of Figure 3.
@@
In the endurance test, ToddlerBot starts with a fully charged battery, running the walking RL policy while stepping in place. ToddlerBot achieved the longest streak of 19 minutes without falling. Over time, increased motor temperatures gradually pushed it outside the policy’s training distribution, leading to more frequent falls. ToddlerBot withstands up to 7 falls before breaking, but even then, repairs are quick—requiring only 21 minutes of 3D printing and 14 minutes of assembly, including removing the damaged part, installing the replacement, and performing zero-point calibration.
@@
Capability: Push-ups and Pull-ups. To demonstrate expressive and dynamic motions, we program push-ups and pull-ups in our keyframe software and perform zero-shot sim-to-real transfer (Figure 4). For pull-ups, we use an AprilTag to help ToddlerBot locate the horizontal bar. Both tasks require strong limbs, balanced upper-lower body strength, and precise coordination, particularly when ToddlerBot transitions from a planking pose to standing after push-ups and when it releases the horizontal bar and lands after pull-ups. These open-loop transfers require only a single motion trajectory designed in simulation, highlighting the fidelity of our digital twin.
Figure 4: Experiment Results. We present four different tasks: push-up, pull-up, bimanual, and full-body manipulation, showing ToddlerBot’s capability in challenging loco-manipulation tasks.
@@
Figure 5: Long-horizon Collaboration. In this task, two instances of ToddlerBot, Arya and Toddy, collaborate to clean up a toy session. (1) The task begins with a pink octopus on the table and a purple octopus on the ground. (2) Arya picks up the pink octopus from the table and places it in the wagon. (3) Arya walks to the wagon handle. (4) Arya grasps the handle while Toddy walks over. (5) Arya pushes the wagon toward the purple octopus. (6) Toddy reaches the pickup position. (7) Toddy kneels and picks up the purple octopus. (8) Finally, Arya and Toddy leave side by side.
@@
ML-Compatibility: Omnidirectional Walking. To demonstrate ToddlerBot’s locomotion capabilities, we train RL walking policies to follow a square trajectory with a predefined velocity profile. The right part of Figure 3 presents the results, with real-world tracking data collected via motion capture. Due to RL policy limitations, both simulation and real-world tracking deviate from the command, primarily because the learned walking policy struggles with in-place rotation, causing translation offsets. However, the sim-to-real gap is notably smaller than the tracking gap, supporting a successful zero-shot transfer. Additionally, we report a position tracking error variance of $0.018\leavevmode\nobreak\ \mathrm{m}$, which demonstrates good repeatability (more details in Appendix 8.14).
@@
ML-Compatibility: Vision-based Manipulation. We show ToddlerBot’s ability to perform bimanual manipulation and full-body manipulation by transferring octopus toys from a table and the ground to a wagon. Both tasks are trained with RGB-based diffusion policy [48] with 60 demonstrations. Across 20 test trials, we achieve a 90% success rate for bimanual manipulation and 75% for full-body manipulation. We leverage a combination of open-loop motions and closed-loop policies to enhance data collection efficiency. In the bimanual task, the torso rotating and releasing motions are open-loop, while in full-body manipulation, kneeling down is open-loop. ToddlerBot’s onboard computing runs a 300300M parameter diffusion policy with about $100\leavevmode\nobreak\ \mathrm{ms}$ latency, enabling real-time operation. Results are shown in Figure 4.
@@
ML-Compatibility: Skill Chaining We test ToddlerBot’s ability to combine loco-manipulation skills with wagon pushing. The results are shown in Appendix 8.15. To push the wagon, ToddlerBot first executes a diffusion policy to grasp the handle, while maintaining that pose, switches to the RL policy to walk forward. To enable walking while maintaining the grip, we sample the robot’s end pose from the grasping policy’s training data (60 demonstrations) during RL training.
@@
Reproducibility: Hardware and Policies To demonstrate hardware reproducibility, we recruit a CS-major student with no prior hardware experience to assemble a second ToddlerBot using our open-source assembly manual and videos. The student independently completes the assembly in three days, including the time for 3D printing. The open-source community has also reported successful reproductions of ToddlerBot, and most of them were completed in one week (Appendix 8.16). For policy reproducibility, we run the manipulation policy trained on data collected with one instance on the other instance, achieving the same success rate of 90% across 20 trials. We also successfully transferred the RL walking policy between both robots. To further showcase the equivalent performance of both ToddlerBot instances, we have them collaborate on a long-horizon toy tidy-up session, as shown in Figure 5.
6 Conclusions